Find the square roots of the following numbers by the Prime Factorisation Method. i. 729 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

729 = 3×3×3×3×3×3×1
⇒ 729 = (3×3)×(3×3)×(3×3) ⇒ 729 = (3×3×3)×(3×3×3) ⇒ 729 = (3×3×3)2⇒√729=3×3×3=27
Thanks